# Mechanisms and Computational Design of Multi-Modal End-Effector with Force Sensing using Gated Networks

### v3

@@ Abstract
In limbed robotics, end-effectors must serve dual functions, such as both feet for locomotion and grippers for grasping, which presents design challenges. This paper introduces a multi-modal end-effector capable of transitioning between flat and line foot configurations while providing grasping capabilities. MAGPIE integrates 8-axis force sensing using proposed mechanisms with hall effect sensors, enabling both contact and tactile force measurements. We present a computational design framework for our sensing mechanism that accounts for noise and interference, allowing for desired sensitivity and force ranges and generating ideal inverse models. The hardware implementation of MAGPIE is validated through experiments, demonstrating its capability as a foot and verifying the performance of the sensing mechanisms, ideal models, and gated network-based models.
@@
11footnotetext: Y. Tanaka and D. Hong are with the Department of Mechanical and Aerospace Engineering, UCLA, Los Angeles, CA, USA. 2A. Zhu is with the Department of Computer Science, UCLA, Los Angeles, CA, USA. 3R. Lin and A. Zhu are with the Electrical and Computer Engineering Department, UCLA, Los Angeles, CA, USA. {yusuketanaka, alvister88, richardlin, mehtank, dennishong}@g.ucla.edu. Y. Tanaka and A. Zhu assert joint first authorship.
@@ I Introduction
The field of limbed robotics, which uses its limbs as both arms and legs, has seen significant growth in recent years thanks to their multi-modal capabilities. Limbed robots can interact with objects without a dedicated manipulation arm [1], can traverse various complex terrains [2] and uneven surfaces [3], and can climb over obstacles [4]. These abilities expand beyond traditional legged robots’ applications, such as search and rescue [2] or planetary exploration [5].
However, the multi-modal nature of limbed robotics presents challenges for end-effector designs since they need to act as both feet for locomotion and grippers for manipulation. As a foot, passive and mechanical compliance introduces challenges in legged robot attitude controls since they are not actively controllable [6]. Contact detection and contact force sensing play a significant role in the legged robot’s locomotion [7]. Grasping force measurements enable both limb movement and grasping to be force-controlled, allowing for more compliant motions [8], [9]. A 6-axis force/torque (F/T) sensor is commonly employed on the ankle, such as in [6]. However, compact force sensors are essential when the sensor cannot be positioned on the ankle or when tactile information from the foot is required for a specific task. [10]. Particularly for highly dynamic bipedal robots, narrow and line feet are more suitable since the contact dynamics and modeling are simplified [11]. However, the line foot design has limited space for contact sensors and is inherently unstable in static and quasi-static cases. These challenges motivate a new end-effector design that can realize multiple configurations while achieving high degrees of force sensing.
@@
Hence, we present MAGPIE (Multi-modal Adaptive Gripper for multi-Pedal Impact-resilient End-effector), a two-finger parallel gripper that functions as a multi-modal foot with an integrated 8-axis force sensing. MAGPIE can transition between flat and line foot configurations with grasping capabilities, providing adequate modality for limbed robotics that necessitates contacting and grasping end-effectors. The design of sensing mechanisms, accounting for noise, and generating inverse models. Our main contributions are:
@@
  • Computational design framework [^1] Codes available on https://suke0811.github.io/magpie_sim/ for 3D hall effect sensor-based 8-axis force sensing mechanisms.
@@ II-A Grasping End-Effector as a Foot
Gripping and grasping end-effectors have been developed to improve foot traction on challenging terrains, to grasp, adhere, and climb on objects, and to enable locomotion and manipulation with the same end-effector. Adhesion types of grasping mechanisms such as magnetic [12], suction [13], and gecko [14] are common mechanisms for legs and grasping. Although these adhesive types are designed for relatively flat surfaces, they could be installed on fingers [15]. The granular jamming-based end-effector, conventionally used for grasping objects through its variable compliance [16], has improved the legged robot tractions over various natural terrains [17]. This passive adaptation limits control of end-effector geometry, and adding contact sensors is a non-trivial task. Multi-fingered grippers with spine contact surface [18] can grasp objects or terrains, while the spine needles improve microscale contacts with rough rocky surfaces [19], [20]. However, this type of high degree of freedom (DoF) grippers are less suitable as a foot end-effector for dynamic locomotion and are meant for quasi-static locomotion. A multi-modal gripper can grasp a rock and transform it into a wheel for rover locomotion [21]. The GOAT gripper [22] has two linear flat fingers and has been effective for climbing and quadruped locomotion, but exhibits structural compliance, and the foot is too narrow to stand on statically. Hence, it was not suitable for bipedal locomotion [23].
@@ II-B Force, Contact, and Tactile Sensing on end-effectors
Single and multi-axis F/T sensors are common choices used in legged robots [24], [25]. These sensors typically consist of strain gauges to measure deflections caused by applied forces [8]. However, multi-axis F/T sensors designed for the large force and torque ranges required by adult-sized humanoids are often too large to fit in confined space [10]. Tactile skin sensors, such as arrays of strain gauges or vision cameras [26], are designed for more precise manipulation tasks but tend to be less durable. Acoustic and viscous flow tactile sensors offer durability and a relatively higher range of tactile force sensing [27], though they are limited by the use of audible signals, restricting their use. Multi-modal foot with acoustic, tactile, and capacitive sensors [28] has been effective in obtaining high dimensional sensing with adaptability for an adult-size bipedal robot, though their focus is on rough terrain classification and locomotion. hall effect sensor-based force sensing mechanisms have been explored for more than 2-axis [29] and [30]. However, the multi-axis hall effect sensing mechanism should be designed to generate proper magnetic field changes at the sensor [31], which requires simulation or empirical testing.
MAGPIE overcomes these limitations by employing four three-axis hall effect sensors that detect ground reaction and grasping forces at the edges of each finger, totaling 8-axis force sensing per gripper. Due to magnetic field sensing’s nonlinearity, we introduce a computational design framework to search for various design parameters. This framework also generates an idealized inverse sensing model.
@@ III-C Circuitry and Sensors
The electronics of the MAGPIE gripper, shown in Fig. 5, are designed to be modular and self-contained, allowing for straightforward integration with the rest of the robotic system. The electrical circuit boards are developed using the hardware description language, Polymorphic Blocks [32] and its human-computer interface [33], which enables schematic, component-level design, electrical modeling, and design verification programmatically. The MAGPIE PCB board integrates a BLDC motor, a 6-axis inertial measurement unit (IMU), an RGB camera, a current/voltage sensor, a color touch display, DC-DC converters, and interfaces via UART and CANBus. The IMU and the cameras will benefit the future end-effector state estimation and visual servo. The four 3-axis hall effect sensors are mounted on custom flexible PCBs with stiffers to simplify sensor placement and reduce hardware complexity. This flexible PCB design minimizes wiring and potential points of failure, improving sensor reliability and ease of installation in MAGPIE.
@@ IV-A Flexure Analytical and Simulation Model
Where $EI$ represents the flexural rigidity, $E$ is Young’s modulus, and $I$ is the second moment of the area of the beam. $w(x)$ denotes the vertical deflection of the beam at a position $x$ along its length. $N(x)$ is the axial force in the beam, and $q(x)$ represents the distributed transverse load per unit length.
@@
When the axial force $N(x)$ is negligible, (1) simplifies bending equation as:
$EI\frac{d^{4}w(x)}{dx^{4}}=P\cdot\frac{L-x}{L}$ (3)
The beam edge’s angle is derived from the slope of the deflection curve, $w(x)$. The derivative, $\frac{dw(x)}{dx}$, gives the slope at any point along the beam. Using the expression for the second derivative of $w(x)$, the slope at the edge can be obtained as:
@@ IV-C Magnetic Interference and Uncertainty
The force sensing is sensitive to magnetic interference. The following are the major sources. 1) Earth’s magnetic field, 2) sensor and magnet misalignment, 3) nearby permanent magnet, i.g., another sensor magnet when the gripper is closed, and 4) permanent magnet inhomogeneity.
Conventionally, the external magnetic field interference can be conveyed by having a significantly stronger permanent magnet for the sensor [29]. Here, we incorporate the model to determine the design parameters, changing the mechanism design parameters to mitigate the expected disturbances.
@@ IV-C3 Permanent magnet nearby
In MAGPIE, another permanent magnet can be nearby when the gripper is closed. Although this could be tried to shield or compensate by knowing the gripper finger current distance, our computational framework allows us to design sensing mechanisms that do not affect the measurements more than the desired level at most.
@@ IV-C4 Permanent magnet impurity
Figure 6: Simulated magnetic field changes at the sensor frame. The magnets purely move in the $x$ direction. For axial metric magnets, the motion in $x$ and $y$-axis generates the identical signal.
@@ IV-D1 Mechanical Constraints
Where $M(x)$ is the bending moment at position $x$, $c$ is the distance from the neutral axis to the outermost fiber, and $I$ is the second moment of area. To avoid fatigue failure, the stress amplitude $\sigma_{a}=\frac{\sigma_{\text{max}}-\sigma_{\text{min}}}{2}$ should correspond to an indefinite fatigue life $N_{f}$ according to the material’s S-N curve.
IV-D2 Geometric Constraints
In addition to geometric constraints due to the sensor and magnet sizes, the force-sensing range can be limited by having a collision under excessive forces. The sensor survival force range can be increased mechanically by limiting the maximum bending deflection.
IV-D3 Magnetic Signal Integrity Constraints
The estimated magnetic field strength at the sensor must be within the 3D hall effect sensor range. The possible magnetic disturbances discussed in IV-C introduce uncertainty in the sensing.
@@ IV-E Force Sensing Model
We incorporate real-world data from the MAGPIE hardware to account for uncertainties, such as sensor nonlinearity and hysteresis, not considered in our analytical models. Given the time-dependent nature of hysteresis, we select a stacked Gated Recurrent Unit (GRU) [36], which takes sensor measurements as inputs and outputs estimated mean forces and their uncertainties. Estimating uncertainty is vital to detect unmodeled magnetic field disturbances, such as external magnets, since more significant uncertainty can indicate abnormal operating conditions.
@@ V Results and Hardware Experiments
V-A Magent Shapes and Sizes
@@ V-B Sensitivity, Magnet Parameters with Self-Interference
Here, using the computational design framework, we simulate the relationship among the maximum sensitivity, $\max(\mathcal{S})$ magnet sizes, and the relative distance to the hall effect sensor. The magnet is moved along with the $x$ axis of the sensor in the range of $(-3.0,3.0)$ m, and the simulation results in Fig. 6b. The identical size of the magnet is placed at $1.5$ mm to represent the case where the gripper is closed, i.e. the closest distance to the other foot magnet.
@@ V-B1 Sensing Mechanisms Design Search
Figure 8: The ground truth, the ideal model, the stacked GRU with 2-axis and 3-axis inputs.
@@ V-B2 Design Selection
Hardware MAGPIE employs A31301 3D linear hall-effect sensors by ALLEGRO Microsystems with a ±$\pm 2000$ Gauss sensing range and $0.1$ Gauss of resolution, which comes with internal compensation such as temperature effects. Through the design search, the beam deflection range was set to $0.5$ mm, and the flexure beam was manufactured using wire EDM. The effective beam length, thickness, width were $30$, $5$ and $15$ mm, respectively Due to the sensor packages and structure limitation, the sensor and the magnet distance $\mathcal{M}_{G}$ had to be at least $1.5$ mm. From Fig. 6c, The magnet $\mathcal{M}_{D}=0.003$ m was selected for the highest sensitivity. The sensitivity is nonlinear for the sensing range, which is complex in the coaxial force cases, which is handled through a stacked GRU framework. When the gripper is closed, this mechanism’s configuration has less than ±$\pm 3$ resolutions of A31301 from the simulation.
@@ V-C GRBF and GRU Model on Hardware
The ground truth force was recorded using Bota Rokubi force/torque sensors attached to each flexure axis at $100$ Hz to determine the relationship between the force applied and the magnetic sensor measurements. Each hall effect sensor has bias calibrated to account for the sensor-magnet misalignment and variances. All four hall effect sensors run at $1000$ Hz on board. The GRBF ideal model force estimate is from the simulation in Section. IV-E, which does not incorporate the axis coupling effects. Fig. 8 compares the ground truth applied forces, idea model estimated forces, and the stacked GRU estimated forces with $2$-axis and 3-axis sensor inputs. Each model estimation root square mean error (RSME) and statistical results are listed in Table. I.
A sinusoidal lateral $F_{x}$ and longitudinal $F_{z}$ force is applied with varying amplitude with an average frequency of $0.5$ Hz as shown in Fig. 8a. The error normalized histogram in Fig. 8b was obtained over the $100$ s test. The GRBF ideal model has shown statistical relationships to the hardware sensor-force correlation but underestimates the force at high amplitude. All models showed no statistically meaningful differences in the phase delay in estimation except for 2-Axis GRU $F_{z}$. The 2-axis GRU model struggled particularly when both axes experienced significant deviation, such as at $5.7$ and $7.7$ s. The 3-axis GRU has handled the coupled cases better than the 2-axis GRU with the same amount of data due to additional redundant information indicating the coupling of both axes. The uncertainty estimation increases as the 3-axis GRU deviates significantly from the ground truth, but the uncertainty estimations are low, around zero, regardless of the degree of deviation. Non-uniform normalization may be necessary to improve the variance estimation around zero to force learning in these regions [37].

### v4

@@ Abstract
In limbed robotics, end-effectors must serve dual functions, such as both feet for locomotion and grippers for grasping, which presents design challenges. This paper introduces a multi-modal end-effector capable of transitioning between flat and line foot configurations while providing grasping capabilities. MAGPIE integrates eight-axis force sensing using proposed mechanisms with Hall effect sensors, enabling both contact and tactile force measurements. We present a computational design framework for our sensing mechanism that accounts for noise and interference, allowing for desired sensitivity and force ranges and generating ideal inverse models. The hardware implementation of MAGPIE is validated through experiments, demonstrating its capability as a foot and verifying the performance of the sensing mechanisms, ideal models, and gated network-based models.
@@
11footnotetext: 1Y. Tanaka and D. Hong are with the Department of Mechanical and Aerospace Engineering, UCLA, Los Angeles, CA, USA. 2A. Zhu is with the Department of Computer Science, UCLA, Los Angeles, CA, USA. 3R. Lin, A. Mehta, and A. Zhu are with the Electrical and Computer Engineering Department, UCLA, Los Angeles, CA, USA. {yusuketanaka, alvister88, richardlin, mehtank, dennishong}@g.ucla.edu. Y. Tanaka and A. Zhu assert joint first authorship.
@@ I Introduction
The field of limbed robotics, which uses its limbs as both arms and legs, has seen significant growth in recent years thanks to its multi-modal capabilities. Limbed robots can interact with objects without a dedicated manipulation arm [1], can traverse various complex terrains [2] and uneven surfaces [3], and can climb over obstacles [4]. These abilities expand beyond traditional legged robots’ applications, such as search and rescue [2] or planetary exploration [5].
However, the multi-modal nature of limbed robotics presents challenges for end-effector designs since they need to act as both feet for locomotion and grippers for manipulation. As a foot, passive and mechanical compliance introduces challenges in legged robot attitude controls since they are not actively controllable [6]. Contact detection and contact force sensing play a significant role in the legged robot’s locomotion [7]. Grasping force measurements enable both limb movement and grasping to be force-controlled, allowing for more compliant motions [8], [9]. A 6-axis force/torque (F/T) sensor is commonly employed on the ankle, such as in [6]. However, compact force sensors are essential when the sensor cannot be positioned on the ankle or when tactile information from the foot is required for a specific task [10]. The more dynamic capable actuator designs [11], [12] allow narrow and line feet configurations, which is more suitable for dynamic bipedal robots since the contact model is simplified [13]. However, the line foot design has limited space for contact sensors and is inherently unstable in static and quasi-static cases. These challenges motivate a new end-effector design that can realize multiple configurations while achieving high degrees of force sensing.
@@
Hence, we present MAGPIE (Multi-modal Adaptive Gripper for multi-Pedal Impact-resilient End-effector), a two-finger parallel gripper that functions as a multi-modal foot with integrated eight-axis force sensing. MAGPIE can transition between flat and line foot configurations with grasping capabilities, providing adequate modality for limbed robotics that necessitates contacting and grasping end-effectors. Our main contributions are:
@@
  • Computational design framework [^1] Codes available on https://suke0811.github.io/magpie_sim/ for 3D Hall effect sensor-based eight-axis force sensing mechanisms.
@@ II-A Grasping End-Effector as a Foot
Gripping and grasping end-effectors have been developed to improve foot traction on challenging terrains, to grasp, adhere, and climb on objects, and to enable locomotion and manipulation with the same end-effector. Adhesion types of grasping mechanisms such as magnetic [14], suction [15], and gecko [16] are common mechanisms for legs and grasping. Although these adhesive types are designed for relatively flat surfaces, they could be installed on fingers [17]. The granular jamming-based end-effector, conventionally used for grasping objects through its variable compliance [18], has improved the legged robot tractions over various natural terrains [19]. This passive adaptation limits control of end-effector geometry, and adding contact sensors is a non-trivial task. Multi-fingered grippers with spine [20] can grasp objects or terrains, while the spine needles improve microscale contacts with rough surfaces [21], [22]. However, this type of high degree of freedom (DoF) grippers are less suitable as a foot for dynamic locomotion and are meant for quasi-static locomotion. A multi-modal gripper can grasp a rock and transform it into a wheel [23]. The GOAT gripper [24] has two linear flat fingers and has been effective for climbing and quadruped locomotion, but exhibits structural compliance, and the foot is too narrow to stand stationary. Hence, it was not suitable for bipedal locomotion [25].
@@ II-B Force, Contact, and Tactile Sensing on end-effectors
Single and multi-axis F/T sensors are common choices in legged robots [26], [27]. These sensors typically consist of strain gauges to measure deflections caused by applied forces [8]. However, multi-axis F/T sensors designed for the large force and torque ranges required by adult-sized humanoids are often too large to fit in confined space [10]. Tactile skin sensors, such as arrays of strain gauges or vision cameras [28], are designed for more precise manipulation tasks but tend to be less durable. Acoustic and viscous flow tactile sensors offer durability and a relatively higher range of force sensing [29], though the use of audible signals restricts their use case. Multi-modal foot with acoustic, tactile, and capacitive sensors [30] has been effective in obtaining high dimensional sensing with adaptability for an adult-size bipedal robot, though their focus is on rough terrain classification and locomotion. Hall effect sensor-based force sensing mechanisms have been explored for more than 2-axis [31], [32]. However, the multi-axis Hall effect is sensitive and nonlinear [33], which requires simulation or empirical tests.
MAGPIE overcomes these limitations by employing four three-axis Hall effect sensors that detect ground reaction and grasping forces at the edges of each finger, totaling eight-axis force sensing per gripper. Due to magnetic field sensing’s nonlinearity, we introduce a computational design framework to search for various design parameters. This framework also generates an idealized inverse sensing model.
@@ III-C Circuitry and Sensors
The electronics of the MAGPIE gripper, shown in Fig. 5, are designed to be modular and self-contained, allowing for straightforward integration with the rest of the robotic system. The electrical circuit boards are developed using Polymorphic Blocks [34] and its human-computer interface [35], which enables schematic, component-level design, electrical modeling, and design verification. The MAGPIE PCB board integrates a BLDC motor, a 6-axis inertial measurement unit (IMU), an RGB camera, a current/voltage sensor, a color touch display, DC-DC converters, and interfaces via UART and CANBus. The IMU and the cameras will benefit the future end-effector state estimation and visual servo. The 3-axis Hall effect sensor is soldered on a custom flexible PCB to minimize wiring and simplify sensor placement, reducing potential points of failure.
@@ IV-A Flexure Analytical and Simulation Model
Where $E$ is Young’s modulus, and $I$ is the second moment of the area of the beam. $w(x)$ denotes the vertical beam deflection at a position $x$ along its length. $N(x)$ is the beam axial force, and $q(x)$ represents the distributed transverse load per unit length.
@@
The derivative, $\frac{dw(x)}{dx}$, gives the slope at any point along the beam. Using the expression for the second derivative of $w(x)$, the slope at the edge can be obtained as:
@@ IV-C Magnetic Interference and Uncertainty
The Hall effect sensor-based force sensing is sensitive to magnetic interference. Conventionally, a strong permanent magnet is simply employed to convey such interference [31]. Here, we incorporate interference models to determine the design parameters, changing the mechanism design parameters to mitigate the expected disturbances.
@@ IV-C3 Permanent magnet nearby
In MAGPIE, another permanent magnet can be nearby when the gripper is closed. Instead of shielding or compensating for this interference, our computational framework allows us to design sensing mechanisms that do not affect the measurements more than the desired level at most.
@@ IV-C4 Permanent magnet impurity
Figure 6: Simulated magnetic field changes at the sensor frame. The magnets purely move in the $x$ direction. For axial metric magnets, the motion in the $x$ and $y$-axis generates the identical signal.
@@ IV-D Constraints
Where $M(x)$ is the bending moment at position $x$, and $c$ is the distance from the neutral axis to the outermost fiber. To avoid fatigue failure, the stress amplitude $\sigma_{a}=\frac{\sigma_{\text{max}}-\sigma_{\text{min}}}{2}$ should correspond to an indefinite fatigue life $N_{f}$ according to the material’s S-N curve.
In addition to geometric constraints due to the sensor and magnet sizes, the force-sensing range can be limited by having a collision under excessive forces. The sensor survival force range can be increased mechanically by limiting the maximum bending deflection. The estimated magnetic field strength at the sensor must be within the 3D Hall effect sensor range. The possible magnetic disturbances discussed in IV-C should be considered as well.
@@ IV-E Force Sensing Model
We incorporate real-world data from the MAGPIE hardware to account for uncertainties, such as sensor nonlinearity and hysteresis, not considered in our analytical models. Given the time-dependent nature of hysteresis, we select a stacked Gated Recurrent Unit (GRU) [38], which takes sensor measurements as inputs and outputs estimated mean forces and their uncertainties. Estimating uncertainty is vital to detect unmodeled magnetic field disturbances, such as external magnets, since significant uncertainty can indicate abnormal operating conditions.
@@ V Results and Hardware Experiments
V-A Magnet Shapes and Sizes
@@ V-B Sensitivity, Magnet Parameters with Self-Interference
Here, using the computational design framework, we simulate the relationship among the maximum sensitivity, $\max(\mathcal{S})$ magnet sizes, and the relative distance to the Hall effect sensor. The magnet is moved along with the $x$ axis of the sensor in the range of $(-3.0,3.0)$ mm, and the simulation results in Fig. 6b. The identical size of the magnet is placed at $1.5$ mm to represent the case where the gripper is closed, i.e., the closest distance to the other foot magnet.
@@ V-B1 Sensing Mechanisms Design Search
Figure 8: The ground truth, the ideal model, and the stacked GRU with 2-axis and 3-axis inputs results from hardware.
Figure 9: Model error distribution for the data in Fig. 8.
@@ V-B2 Design Selection
Hardware MAGPIE employs A31301 3D linear hall-effect sensors by ALLEGRO Microsystems with a ±$\pm 2000$ Gauss range and $0.1$ Gauss of resolution, which comes with internal compensation such as temperature effects. The design search set the beam deflection range to $0.5$ mm, and the flexure beam was manufactured using wire EDM. The effective beam length, thickness, width were $30$, $5$ and $15$ mm, respectively Due to the sensor packages and structure limitation, the sensor and the magnet distance $\mathcal{M}_{G}$ had to be at least $1.5$ mm. From Fig. 6c, The magnet $\mathcal{M}_{D}=0.003$ m was selected for the highest sensitivity. The sensitivity is nonlinear for the sensing range, which is complex in the coaxial force cases and is handled through a stacked GRU framework. When the gripper is closed, this mechanism’s configuration has less than ±$\pm 3$ resolutions of A31301 from the simulation.
@@ V-C GRBF and GRU Model on Hardware
V-C1 Data Collection
The ground truth force was recorded with two Bota Rokubi F/T sensors attached to each flexure axis at $100$ Hz to determine the relationship between the force applied and the magnetic sensor measurements. Each F/T and Hall effect sensor was bias-calibrated to account for the misalignment and variances of the magnet. All four Hall effect sensors run at $1000$ Hz on board. The GRBF ideal model force estimate is from the simulation in Section. IV-E. Fig. 8 compares the ground truth applied forces, idea model estimated forces, and the stacked GRU estimated forces with $2$-axis and 3-axis sensor inputs. Each model estimation root square mean error (RSME) and statistical results are listed in Table. I. A sinusoidal lateral $F_{x}$ and longitudinal $F_{z}$ force is applied with varying amplitude with an average frequency of $0.5$ Hz as shown in Fig. 8. The error normalized histogram in Fig. 9 was obtained over the $100$ s test.
V-C2 Statistical Analysis and discussion
The GRBF ideal model has shown statistical relationships to the hardware sensor-force correlation but underestimates the force at high amplitude. All models showed no statistically meaningful differences in the phase delay in estimation except for 2-Axis GRU $F_{z}$. The 2-axis GRU model struggled particularly when both axes experienced significant deviation, such as at $5.7$ and $7.7$ s. The 3-axis GRU has handled the coupled cases better than the 2-axis GRU with the same amount of data due to additional redundant information indicating the coupling of both axes. The uncertainty estimation increases as the 3-axis GRU deviates from the ground truth.
The 3-axis GRU model outperforms the ideal model, the 2-axis GRU model. However, the ideal model estimation was consistently closer to the ground truth around $<25$ N or $<10$ of the force range. This is potential because the lower force range dataset is more sensitive to the F/T and Hall effect sensors’ initial calibration. In such cases, GRU requires more data to learn biases. In the future, adding the ideal model estimation as GRU input can enhance the training data efficiency and improve the estimation quality since the ideal model works as a warm start initial guess. The uncertainty estimation around zero can be potentially improved with non-uniform normalization to force learning in these regions [39].
V-C3 Computation Cost and External Interference